OpenStudy (anonymous):

How many liters of a 60% acid solution must be added to 14 liters of a 10% acid solution to produce a 25% acid solution?

OpenStudy (amistre64):

i loathe word problems :)

OpenStudy (amistre64):

x(.6) + 14(.1) = (x+14)(.25) .6x + 1.4 = .25x + 3.5 .6x - .25x = 3.5 - 1.4 x(.6 - .25) = 2.1 x(.35) = 2.1 x= (2.1)/(.35) = 210/35 x = 6 Double check: 6(.6) + 14(.1) = 20(.25) 3.6 + 1.4 = 5 5=5 (its good) :)